OpenStudy (anonymous):

What are the steps to solve (1)/(2)+(1)/(3)=(1)/(x) does x=5/6?

OpenStudy (anonymous):

1/2+ 1/3 = 1/x FIRST TAKE L.C.M of 2 and 3

OpenStudy (anonymous):

Actually, 1/x = 5/6

OpenStudy (anonymous):

Now, can u find x

OpenStudy (anonymous):

Thank you! I understand now :)

OpenStudy (anonymous):

yw